# Practical X-ray Gastric Cancer Screening Using Refined Stochastic Data Augmentation and Hard Boundary Box Training

### v4

@@ Abstract
We propose an unprecedented gastric cancer diagnostic support system for gastric X-rays, which can be performed by technicians, allowing more people to be screened compared to endoscopy, the current mainstream method that can only be performed by physicians. The system is based on a general deep learning-based object detection model and introduces two novel techniques: refined stochastic gastric image augmentation (R-sGAIA) and hard boundary box training (HBBT).
A R-sGAIA is a probabilistic gastric fold region enhancement method that provides more learning patterns for cancer detection models.
A HBBT is an efficient and versatile training method for common object detection models, enabling the use of unannotated negative (i.e., healthy control) samples that cannot be used in conventional detection models. This reduces false positives and improves overall performance.
The proposed system achieves a sensitivity of 90.2% for gastric cancer, exceeding that of an expert (85.5%), with 42.5% of the detected candidate box regions showing cancerous lesions and a processing time of 0.51 seconds per image. It also achieves a 5.9-point higher F1 score compared to methods using the same object detection model and state-of-the-art data augmentation. The system efficiently guides radiologists on where to focus, significantly reducing their workload.
@@ 1 Introduction
Gastric cancer is the third most deadly cancer globally, with a poor prognosis, especially in advanced stages (Rawla and Barsouk, 2019). However, early detection before metastasis can significantly improve outcomes, making early diagnosis and appropriate treatment crucial. Gastric cancer is primarily diagnosed through endoscopy or radiography. Endoscopy reports a sensitivity (SE) of 95.4%, which is superior to other methods, such as gastric radiography or X-ray, at 85.5% (Hamashima et al., 2013). However, endoscopy can only be performed by a physician, and the time and cost involved limit the number of people who can be screened. In contrast, gastric X-rays can be performed by radiographers, generating a large number of images, which makes this method suitable for mass screening – a significant advantage over endoscopic diagnosis. In Japan, a mass screening system was established, and in 2019, at least 3.87 million people underwent gastric X-rays for gastric cancer screening. Of these, 183,000 people were indicated for further testing, and 2,553 cases of gastric cancer were ultimately detected. By comparison, only 374,000 people underwent endoscopy for the same purpose (Matsuura, 2023). Although endoscopy’s effectiveness has led to a shift toward this diagnostic method, developing superior diagnostic support technology for gastric radiography would significantly contribute to the early detection and reduction of mortality from gastric cancer.
@@
Gastric X-ray images are later interpreted by doctors to diagnose the stomach’s shape and mucosal atrophy (Uemura et al., 2001; Roder, 2002; Correa et al., 2004). An additional advantage of X-rays is their ability to reveal minute irregularities indicative of stage IIc lesions (early-stage tumors with slight indentations) – a subtle elevation within the indentation that lies beneath it. However, accurately reading these images requires a high level of skill, which many gastrointestinal radiologists may lack. When performed by a skilled physician, there is a high likelihood of assessing the extent of cancerous invasion deep within the surface of an early-stage cancer based on changes in the microscopic structure of the surface. The limitations of this diagnostic method lie in its dependence on extensive physician experience and the particular difficulty in detecting early-stage cancers, leading to a lower cancer detection capacity than that of endoscopy (Hibino et al., 2023). Therefore, there is a pressing need for an automated screening assistance system that can accurately and promptly detect gastric abnormalities from gastric X-ray images in clinical settings.
@@
Healthy gastric folds are thin, smooth, and parallel; however, their thickness and size can be altered by gastric diseases such as inflammation, infection, ulcers, and tumors, leading to irregular bends, depressions, and tears in the lining mucosa (Correa et al., 2004). Most research on diagnostic support using gastric X-rays has focused on detecting gastritis, the precursor to gastric cancer, or the presence of Helicobacter pylori infection, which is believed to be a major cause of gastric cancer (Cooper et al., 1990; Kita, 1992; Minemoto et al., 2010; Nagano and Matsuo, 2010; Abe et al., 2014; Ishihara et al., 2015, 2017; Togo et al., 2019; Kanai et al., 2019, 2020). Among these studies, methods based on deep learning, particularly convolutional neural networks (CNNs), have shown excellent results in detecting gastritis. The main advantage of these deep learning techniques is their ability to automatically capture efficient features for the target task, which is promising in areas where disease features are difficult to define. However, studies directly detecting gastric cancer from radiographs are limited, despite the significance of gastric cancer diagnosis in terms of life expectancy. This may be because gastric cancer, especially in its early stages, is more difficult to diagnose compared to gastritis, where lesions often expand and develop to some degree. Reportedly, even with the latest machine learning techniques, distinguishing between gastritis and gastric cancer on endoscopic images—where signs of the disease are more visible – remains challenging (Horiuchi et al., 2020).
@@
In the field of computer vision, CNN-based object detection methods (e.g., Faster R-CNN (Ren et al., 2015), YOLOv3 (Redmon and Farhadi, 2018), EfficientDet (Tan et al., 2020), and YOLOv7 (Wang et al., 2023)), which can simultaneously detect and classify the location of multiple objects in images, have been proposed and widely used. These methods are particularly valuable in medical applications because they not only detect regions of interest but also provide high explanatory power for the predicted results, as opposed to simply classifying images as benign or malignant (Hirasawa et al., 2018; Li et al., 2018; Okamoto et al., 2019; Laddha et al., 2019; Zhang et al., 2019). For instance, in a study on gastric X-ray images, Laddha et al. (2019) used YOLOv3 to automatically detect gastric polyps, achieving a mean average precision of 0.916. However, these object detection methods require precise annotation of lesion locations prior to training, which is extremely costly. To avoid high annotation costs while still benefiting from the explanatory advantages of object detection models, previous studies on gastritis detection (Ishihara et al., 2017; Togo et al., 2019; Kanai et al., 2019, 2020) used patch-based CNNs, which only required image-level annotations. In the context of gastric cancer, where diagnostic challenges and outcomes are more severe, it is critical that physicians are able to interpret the results provided by diagnostic support systems.
@@
With this motivation, we proposed a diagnostic support system that directly targets gastric cancer using X-ray images, which have rarely been studied, and offers high explanatory power for the results (Okamoto et al., 2019). Thanks to the proposed stochastic gastric image augmentation (sGAIA) based on medical knowledge, this system accurately represented candidate regions of gastric cancer within bounding boxes. In an evaluation using images from patients not included in the training set, the recall (SE) and precision of the detected bounding boxes were 92.3% and 32.4%, respectively. This recall is about 7% higher than that achieved by doctors (85.5%), and a box-by-box evaluation confirmed a true-to-false detection ratio of 1:3, which is sufficient for practical use. While the object detection model trained with sGAIA has reached a practical level of accuracy, further performance improvements are desirable. Currently, widely used object detection methods have the disadvantage of not being able to explicitly use negative (i.e., healthy control) data for training, in addition to the high cost of annotation. When such models are used for lesion detection, only images with lesions are used for training, while control images without lesions are excluded. This limitation presents an opportunity for improvement, which is a key focus of this study. In medical imaging, data are expensive, and we believe that actively using healthy control images for training can enhance the model’s overall ability to detect lesions, effectively reducing annotation costs while maintaining performance. In this study, we proposed a practical gastric cancer screening system that includes two new techniques: refined stochastic gastric image augmentation (R-sGAIA), an improvement of sGAIA (Okamoto et al., 2019), and hard boundary box training (HBBT), which enables the use of negative or healthy control images for training, a feature not previously utilized in object detection neural networks. To suppress high-confidence false positives predicted by the object detector, HBBT registers each of these boxes as a “hard-sample” class and retrains the model until predetermined convergence conditions are met. We investigated 4,724 gastric X-ray images obtained from 145 patients in a clinical setting and confirmed the practicality and effectiveness of the proposed system by evaluating gastric cancer detection performance using a subject-based five-fold cross-validation strategy.
@@
    We proposed an unprecedented gastric cancer diagnostic support system for gastric X-rays, which can be taken by technicians. As a result, this allows more people to be screened, rather than endoscopy, which is currently the mainstream but can only be performed by physicians. The system is based on a general deep learning-based object detection model and includes two novel technical proposals: R-sGAIA and HBBT.
@@
    HBBT is an efficient and versatile training method for common object detection models that allows the use of unannotated negative (i.e., healthy control) samples that otherwise cannot be used for training in conventional detection models. This reduces false positives and thus improves overall performance.
@@
    The SE of the proposed system for gastric cancer (90.2%) exceeds that of the expert (85.5%), while 42.5% of the detected candidate box regions show cancerous lesions, with a processing time of 0.51 s/image. It is 5.9 points higher on the F1 score compared to methods using the same object detection model and state-of-the-art data augmentation. In short, the system efficiently shows the radiologist where to look, greatly reducing the radiologist’s workload.
@@ 2.1 Diagnostic support using gastric X-ray images
In the 1990s, methods were proposed for effectively extracting the pattern of gastric folds by applying binarization as a preprocessing step and estimating the gastric region based on the position of the barium reservoir (Cooper et al., 1990; Kita, 1992). In the 2010s, more empirical studies began to focus on gastric folds (Minemoto et al., 2010; Nagano and Matsuo, 2010; Abe et al., 2014; Ishihara et al., 2015). Ishihara et al. (2015) calculated numerous statistical features (7,760 per case) for H. pylori infection, a major cause of gastric cancer, by analyzing post-infection mucosal patterns and gastric folds. Their support vector machine, trained on data from 2,100 patients (with eight images per person), achieved a SE of 89.5% and a SP of 89.6%.
More recently, deep learning techniques have been applied to the detection of gastritis using patch-based CNNs on gastric X-ray images (Ishihara et al., 2017; Togo et al., 2019; Kanai et al., 2019, 2020). Although these methods targeted gastritis, which is easier to diagnose than gastric cancer, they achieved high detection accuracy and were considered practical for diagnostic assistance. Kanai et al. (2020) used a patch-based CNN to detect H. pylori-associated chronic atrophic gastritis. They trained image patches from both the inside and outside of the stomach and made a diagnosis by taking a majority vote of each estimated result in the stomach region. Manual annotation of regions of interest (ROIs) prior to training, along with self-learning to increase the number of training patches, ultimately yielded a harmonic mean SE and SP of 95.5% for the test data from different patients. While excellent screening results have been achieved for gastritis, the detection of more serious and difficult-to-diagnose gastric cancers has not been well studied. In the case of gastric cancer, it is also crucial to provide an explanation for the results. The proposed method significantly improves upon our previous work on a gastric cancer detection system that can present explainable results (Okamoto et al., 2019), incorporating two new technical advancements.
@@ 2.2 Diagnostic support research for endoscopic images
Wu et al. (2019) used classic VGG16 and ResNet-50 models, achieving a SE of 94.0% and a SP of 91.0% based on 24,549 images for diagnosing early gastric cancer. Li et al. (2020) employed an Inception-v3 model, achieving an SE of 91.2% and an SP of 90.6% based on 2,429 images of gastric magnification endoscopy with narrow-band imaging. Horiuchi et al. (2020) trained a GoogLeNet model with 1,499 images of gastric cancer and 1,078 images of gastritis, achieving an SE of 95.4%, an SP of 71.0%, and an area under the receiver operating characteristic (ROC) curve (AUC) of 0.852. Gong et al. (2023) developed a Convolution and Relative Self-Attention Parallel Network using 3,576 endoscopic images from 205 patients, reporting an F1 score of 0.948 in their experiments with a 7:3 split of training and test images. While these methods produce excellent results, none of them explicitly state whether the training dataset is completely separate from the patients whose images are in the evaluation set. This raises the possibility that images from the same patients were included in both the training and test data, potentially leading to inflated performance scores due to the close similarity between the two datasets.
@@
In contrast, Kanayama et al. (2019) constructed a diagnostic support system using 129,692 gastrointestinal endoscopy images with sophisticated CNN-based networks, including one synthesizer network and two different types of discriminator networks. Importantly, they ensured that images from the same patient were not divided between the training and evaluation sets, thus evaluating performance appropriately. They reported an average precision (AP) of 0.596 for cancer detection. Although the numerical performance of this method may seem lower, it is based on a newer and more sophisticated machine learning model than those used in previously reported methods, representing a more accurate diagnostic capability. They further enhanced their model by generating 20,000 high-resolution lesion images from 1,315 narrow-field-of-view images using their own generative adversarial network (Goodfellow et al., 2014) and adding these to the training set, improving the AP in cancer detection to 0.632. Shibata et al. (2020) constructed and rigorously evaluated a diagnostic model using Mask R-CNN on 1,208 images from 42 healthy subjects and 533 images from 93 unhealthy subjects (including gastric cancer). Their model achieved an SE of 96% for all tumor types, a false positive rate of 0.1/image, and an F1 score of 0.71 for gastric cancer. Ahmad et al. (2023) trained and evaluated a larger dataset than in previous studies (27,200 training and 6,800 testing images) with a novel approach that added Squeeze and Excitation attention blocks to YOLOv7 (Wang et al., 2023), achieving an F1 score of 0.71 (precision = 72%, recall = 69%).
While it is unclear whether proper data partitioning was employed, it is important to note that the number of images used was significantly larger than in other studies. Additionally, the use of more sophisticated methods yielded results different from earlier studies (which reported F1 scores around 90% or higher) where substantial data leakage was suspected. Given that the results were comparable to other reports (Shibata et al., 2020) with clearly distinct data partitions, it can be inferred that the data division was likely appropriate.

### v5

# Practical X-ray Gastric Cancer Diagnostic Support Using Refined Stochastic Data Augmentation and Hard Boundary Box Training

@@ Abstract
A novel gastric cancer detection system using X-rays enables scalable screening.
R-sGAIA enhances gastric folds, providing diverse patterns for cancer detection.
HBBT enables training with unannotated data, reducing false positives effectively.
The system outperforms experts with 90.2% sensitivity, diagnosing at 0.51s per image.
Our R-sGAIA and HBBT are available at https://github.com/IyatomiLab/RsGAIA_HBBT/.
@@ 1 Introduction
Gastric cancer is the third most deadly cancer globally and has a poor prognosis, especially in advanced stages (Rawla and Barsouk, 2019). However, early detection before metastasis can significantly improve outcomes, which makes early diagnosis and appropriate treatment crucial. Gastric cancer is primarily diagnosed through endoscopy or radiography. Endoscopy results in a sensitivity (SE) of 95.4%, which is superior to other methods, such as gastric radiography or X-ray, at 85.5% (Hamashima et al., 2013). However, endoscopy can only be performed by a physician, and the time and cost involved in the procedure limit the number of patients who can be screened. In contrast, gastric X-rays can be taken by radiographers, allowing them to examine many patients in a shorter time. This makes the method suitable for mass screening and provides a significant advantage over endoscopic diagnosis. In Japan, a mass screening system was established, and, in 2019, at least 3.87 million people underwent gastric X-rays for gastric cancer screening. Of these, 183,000 people were indicated for further testing, and 2,553 cases of gastric cancer were ultimately detected. By comparison, only 374,000 people underwent endoscopy for the same purpose (Matsuura, 2023). Although endoscopy’s effectiveness has led to a shift toward this diagnostic method, developing superior diagnostic support technology for gastric radiography would significantly contribute to the early detection and reduction of mortality from gastric cancer.
@@
Gastric X-ray images are later interpreted by doctors to diagnose the stomach’s shape and mucosal atrophy (Uemura et al., 2001; Roder, 2002; Correa et al., 2004). An additional advantage of X-rays over endoscopy is their ability to reveal minute irregularities indicative of stage IIc lesions (early-stage tumors with slight indentations), which appear as a subtle elevation within the indentation that lies beneath it. However, accurately reading these images requires a high level of skill, which many gastrointestinal radiologists may lack. When performed by a skilled physician, there is a high likelihood of assessing the extent of cancerous invasion deep within the surface of an early-stage cancer based on changes in the microscopic structure of the surface. The limitations of this diagnostic method lie in its dependence on extensive physician experience and the particular difficulty of detecting early-stage cancers, which leads to a lower cancer detection capacity than that of endoscopy (Hibino et al., 2023). Therefore, there is a pressing need for an automated diagnosis assistance system that can accurately and promptly detect gastric abnormalities from gastric X-ray images in clinical settings.
@@
Healthy gastric folds are thin, smooth, and parallel; however, their thickness and size can be altered by gastric conditions such as inflammation, infection, ulcers, and tumors that lead to irregular bends, depressions, and tears in the lining mucosa (Correa et al., 2004). Most research on diagnostic support using gastric X-rays has focused on detecting gastritis, the precursor to gastric cancer, or the presence of Helicobacter pylori infection, which is believed to be a major cause of gastric cancer (Cooper et al., 1990; Kita, 1992; Minemoto et al., 2010; Nagano and Matsuo, 2010; Abe et al., 2014; Ishihara et al., 2015, 2017; Togo et al., 2019; Kanai et al., 2019, 2020). Among these studies, methods based on deep learning, particularly convolutional neural networks (CNNs), have shown excellent results in detecting gastritis. The main advantage of these deep learning techniques is their ability to automatically capture efficient features for the target task, which is promising in areas where disease features are difficult to define. However, studies directly detecting gastric cancer from radiographs are limited, despite the significance of gastric cancer diagnosis in terms of life expectancy. This may be because gastric cancer, especially in its early stages, is more difficult to diagnose compared to gastritis, where lesions often expand and develop to some degree. Reportedly, even with the latest machine learning techniques, distinguishing between gastritis and gastric cancer on endoscopic images—where signs of the disease are more visible–remains challenging (Horiuchi et al., 2020).
@@
In the field of computer vision, CNN-based object detection methods (e.g., Faster R-CNN (Ren et al., 2015), YOLOv3 (Redmon and Farhadi, 2018), EfficientDet (Tan et al., 2020), and YOLOv10 (yolov10)), which can simultaneously detect and classify the location of multiple objects in images, have been proposed and widely used. These methods are particularly valuable in medical applications because they not only detect regions of interest but also provide high explanatory power for the predicted results, as opposed to simply classifying images as benign or malignant (Hirasawa et al., 2018; Li et al., 2018; Okamoto et al., 2019; Laddha et al., 2019; Zhang et al., 2019). For instance, in a study on gastric X-ray images, Laddha et al. (2019) used YOLOv3 to automatically detect gastric polyps and achieved a mean average precision of 0.916. However, these object detection methods require precise annotation of lesion locations prior to training, which is extremely costly. To avoid high annotation costs while still benefiting from the explanatory advantages of object detection models, previous studies on gastritis detection (Ishihara et al., 2017; Togo et al., 2019; Kanai et al., 2019, 2020) used patch-based CNNs, which only required image-level annotations. In the context of gastric cancer, where diagnostic challenges and outcomes are more severe, it is critical that physicians are able to interpret the results provided by diagnostic support systems.
With this motivation, we developed a diagnostic support system that directly targets gastric cancer using X-ray images,which have received little attention in the literature, and offers high explanatory power for the results (Okamoto et al., 2019). Thanks to their gastric image augmentation (sGAIA) developed based on medical knowledge, this system accurately represented candidate regions of gastric cancer within bounding boxes. In an evaluation using images from patients not included in the training set, the recall (SE) and precision of the detected bounding boxes were 92.3% and 32.4%, respectively. This recall is about 7% higher than that achieved by doctors (85.5%), and a box-by-box evaluation confirmed a true-to-false detection ratio of 1:3, which is sufficient for practical use.
@@
While the object detection model trained with sGAIA reached a practical level of accuracy, further performance improvements are desirable. Currently, widely used object detection methods have the disadvantage of not being able to explicitly use negative (i.e., healthy control) data for training, in addition to the high cost of annotation. When such models are used for lesion detection, only images with lesions are used for training, while control images without lesions are excluded. This limitation presents an opportunity for improvement, which is a key focus of this study. In medical imaging, data are expensive, and we believe that actively using healthy control images for training can enhance the model’s overall ability to detect lesions, thereby effectively reducing annotation costs while maintaining performance.
Therefore, in this study, we propose a practical gastric cancer diagnostic support system that includes two new techniques: refined stochastic gastric image augmentation (R-sGAIA), an improvement of sGAIA (Okamoto et al., 2019), and hard boundary box training (HBBT), which enables the use of negative or healthy control images for training, a feature not previously utilized in object detection neural networks. To suppress high-confidence false positives predicted by the object detector, HBBT registers each of these boxes as a “hard sample” class and retrains the model until predetermined convergence conditions are met. We investigated 4,724 gastric X-ray images obtained from 145 patients in a clinical setting and confirmed the practicality and effectiveness of the proposed system by evaluating gastric cancer detection performance using a subject-based five-fold cross-validation strategy.
@@
    We propose a novel gastric cancer diagnostic support system for gastric X-rays that can be captured by radiographers. The use of radiographers rather than physicians allows more people to be screened than the current mainstream use of endoscopy, which can only be performed by physicians. The system is based on a general deep learning-based object detection model and includes two novel technical proposals: R-sGAIA and HBBT.
@@
    HBBT is an efficient and versatile training method for common object detection models that allows the use of unannotated negative (i.e., healthy control) samples that otherwise cannot be used for training in conventional detection models. This approach reduces false positives and thus improves overall performance.
@@
    The SE of the proposed system for gastric cancer (90.2%) exceeds that of the expert (85.5%), while 42.5% of the detected candidate box regions show cancerous lesions, with a processing time of 0.51 seconds/image. It is 5.9 points higher on the F1 score compared to methods that that use the same object detection model and state-of-the-art data augmentation. In short, the system efficiently shows the radiologist where to look, thereby greatly reducing the radiologist’s workload.
@@ 2.1 Diagnostic support using gastric X-ray images
In the 1990s, methods were proposed for effectively extracting the pattern of gastric folds by applying binarization as a preprocessing step and estimating the gastric region based on the position of the barium reservoir (Cooper et al., 1990; Kita, 1992). In the 2010s, more empirical studies began to focus on gastric folds (Minemoto et al., 2010; Nagano and Matsuo, 2010; Abe et al., 2014; Ishihara et al., 2015). Ishihara et al. (2015) calculated numerous statistical features (7,760 per case) for H. pylori infection, a major cause of gastric cancer, by analyzing post-infection mucosal patterns and gastric folds. Their support vector machine, trained on data from 2,100 patients (with eight images per person), achieved an SE of 89.5% and a specificity (SP) of 89.6%.
More recently, deep learning techniques have been applied to the detection of gastritis using patch-based CNNs on gastric X-ray images (Ishihara et al., 2017; Togo et al., 2019; Kanai et al., 2019, 2020). Although these methods targeted gastritis, which is easier to diagnose than gastric cancer, they achieved high detection accuracy and were considered practical for diagnostic assistance. Kanai et al. (2020) used a patch-based CNN to detect H. pylori-associated chronic atrophic gastritis. They trained image patches from both the inside and outside of the stomach and made a diagnosis by taking a majority vote of each estimated result in the stomach region. Manual annotation of regions of interest (ROIs) prior to training, along with self-learning to increase the number of training patches, ultimately yielded a harmonic mean SE and SP of 95.5% for the test data from different patients. While excellent screening results have been achieved for gastritis, the detection of more serious and difficult-to-diagnose gastric cancers has not been well studied. In the case of gastric cancer, it is also crucial to provide an explanation for the results. The proposed method significantly improves upon our previous work on a gastric cancer detection system that can present explainable results (Okamoto et al., 2019) by incorporating two new technical advancements.
@@ 2.2 Diagnostic support research for endoscopic images
Wu et al. (2019) used classic VGG16 and ResNet-50 models to achieve an SE of 94.0% and an SP of 91.0% based on 24,549 images for diagnosing early gastric cancer. Li et al. (2020) employed an Inception-v3 model to achieve an SE of 91.2% and an SP of 90.6% based on 2,429 images of gastric magnification endoscopy with narrow-band imaging. Horiuchi et al. (2020) trained a GoogLeNet model with 1,499 images of gastric cancer and 1,078 images of gastritis, thereby achieving an SE of 95.4%, an SP of 71.0% as well as an area under the receiver operating characteristic (ROC) curve (AUC) of 0.852. Gong et al. (2023) developed a convolution and relative self-attention parallel network using 3,576 endoscopic images from 205 patients and reported an F1 score of 0.948 in their experiments with a 7:3 split of training and test images. While these methods produced excellent results, none of them explicitly stated whether the training dataset was completely separate from the patients whose images were in the evaluation set. This raises the possibility that images from the same patients were included in both the training and test data, potentially leading to inflated performance scores due to the close similarity between the two datasets.
@@
In contrast, Kanayama et al. (2019) constructed a diagnostic support system using 129,692 gastrointestinal endoscopy images with sophisticated CNN-based networks, including one synthesizer network and two different types of discriminator networks. Importantly, they ensured that images from the same patient were not divided between the training and evaluation sets, thus evaluating performance appropriately. They reported an average precision (AP) of 0.596 for cancer detection. Although the numerical performance of this method may seem lower, it is based on a newer and more sophisticated machine learning model than those used in previously reported methods and represents a more accurate diagnostic capability. They further enhanced their model by generating 20,000 high-resolution lesion images from 1,315 narrow-field-of-view images using their own generative adversarial network (Goodfellow et al., 2014) and, by adding these to the training set, improved the AP in cancer detection to 0.632. Shibata et al. (2020) constructed and rigorously evaluated a diagnostic model using Mask R-CNN on 1,208 images from 42 healthy subjects and 533 images from 93 unhealthy subjects (including gastric cancer). Their model achieved an SE of 96% for all tumor types, a false positive rate of 0.1/image, and an F1 score of 0.71 for gastric cancer. Ahmad et al. (2023) trained and evaluated a larger dataset than in previous studies (27,200 training and 6,800 testing images) with a novel approach that added squeeze and excitation attention blocks to YOLOv7 (Wang et al., 2023), thereby achieving an F1 score of 0.71 (precision = 72%, recall = 69%). While it is unclear whether proper data partitioning was employed for this method, it is important to note that the number of images used was significantly larger than in other studies. Additionally, the use of more sophisticated methods yielded results different from earlier studies (which reported F1 scores around 90% or higher) where substantial data leakage was suspected. Given that the results were comparable to other reports (Shibata et al., 2020) with clearly distinct data partitions, it can be inferred that the data division was likely appropriate.